How many grams of pure silver must a silversmith mix with a 45% silver alloy to produce 200 grams of a 50% alloy?
Answer by checkley77(12844)

x+.45(200-x)=.50*200
x+90-.45x=100
.55x=100-90
.55x=10
x=10/.55
x=18.18 grams of pure siler is used.
Proof:
18.18+.45(200-18.18)=.50*200
18.18+.45*181.82=100
18.18+81.82=100
100=100